OpenStudy (anonymous):

-3 + 8x - 5 = -8 A. -2 B. -1 C. - 3/4 D. 0

HanAkoSolo (jamierox4ev3r):

-3+8x-5=-8 -8+8x=-8 +8 +8 -------- 8x=0 x=0 the answer is D.